OpenStudy (anonymous):

What is the equation of the line described below written in slope-intercept form? the line passing through point (2, 2) and perpendicular to the line whose equation is y = x

OpenStudy (anonymous):

Just like last question Line perpendicular to the given line will have negative reciprocal slope from the given line that it y = x m = 1 slope of perpendicular line = -1 perpendicular line passes through (2,2) so using point slope equation form y - y1 = m(x-x1) y -2 = (-1)(x -2) y -2 = -x +2 y = -x

OpenStudy (anonymous):

I'm just confused because the answers are y = x - 4 y = x + 4 y = -x + 4 Sorry, I'm really bad at these.

OpenStudy (anonymous):

Sorry last line my mistake y - 2 = -x +2 y = -x +4